OpenStudy (anonymous):

solve by elimination x+y=15 x-y=-3

OpenStudy (anonymous):

x+y=15 x-y=-3 2x = 12 /2 /2 x=6 6 + y = 15 -6 -6 y = 9 The answer is (6, 9)

OpenStudy (anonymous):

The answer is (6, 9)

OpenStudy (anonymous):

thanks guys